Can you please help me solve this equation?
5(-2x+1)= -9(x+1)
Thank you so very much!

Answer by nerdybill(7384)

5(-2x+1)= -9(x+1)
Distributing on the left and right side:
-10x+5= -9x-9
5= x-9
14 = x